OpenStudy (anonymous):

If a side-length of an equilateral triangle is 18, which is its area?

OpenStudy (anonymous):

[sqrt(3)/4 ]* 18^2

OpenStudy (anonymous):

thnks

OpenStudy (anonymous):

i still can not find the right answer though

OpenStudy (anonymous):

i got 60.75

OpenStudy (anonymous):

then i got 243 and niether of them are right

OpenStudy (anonymous):

This is the correct formula so just plug in the values

OpenStudy (anonymous):

130.13

OpenStudy (anonymous):

i did

OpenStudy (anonymous):

the altitude = sqrt(18^2-9^2) = 15.588 area = 1/2 base x altitude = 9 * 15.588 = 140.3